jujy15:

Which set of points includes all of the solutions for y = five halves times x plus three halves? line going through negative 1 comma negative 1 and 1 comma 4 (x, five halves times x plus three halves) for all real numbers (−1, −1), (0, 1.5), (1, 4) (x, y) for all real numbers (−3, −6), (−2, −3.5), (0, 1.5), (2, 6.5), (3, 9)

Mooshroom:

For \[y=\frac{ -3 }{ 2 } x+2\] the set of points that include all solutions are \[(x,y)\]such that\[y=\frac{ -3 }{ 2 }x+2\]; then we have\[(x,y)=(x,\frac{ -3 }{ 2 }x+2)\]. What is important to stress here is that\[(x,-\frac{3}{2}x+2)\]represents all of the solutions, while other points that we are given represent one solution (for example, point (2,-1) is just one solution)

Mooshroom:

@mooshroom wrote:
For \[y=\frac{ -3 }{ 2 } x+2\] the set of points that include all solutions are \[(x,y)\]such that\[y=\frac{ -3 }{ 2 }x+2\]; then we have\[(x,y)=(x,\frac{ -3 }{ 2 }x+2)\]. What is important to stress here is that\[(x,-\frac{3}{2}x+2)\]represents all of the solutions, while other points that we are given represent one solution (for example, point (2,-1) is just one solution)
does this help?